A machine is 12% efficient. If useful work done by the machine is 6 joules then calculate the total amount of heat
1
Expert's answer
2019-03-25T08:53:17-0400

"E = A+Q;\\
A=6~\text{J};\\
k = \frac{A}{E} = 0.12;\\
Q = E-A = \frac{A}{k}-A = \frac{1-k}{k}A = 44~\text{J}."